# Position: Levels of AGI for Operationalizing Progress on the Path to AGI

### v4

@@ 3 Defining AGI: Six Principles
6. Focus on the Path to AGI, not a Single Endpoint. Much as the adoption of a standard set of Levels of Driving Automation (SAE International, 2021) allowed for clear discussions of policy and progress relating to autonomous vehicles, we posit there is value in defining “Levels of AGI.” As we discuss in Section 5 and Section 6, we intend for each level of AGI to be associated with a clear set of metrics/benchmarks, as well as identified risks introduced at each level, and resultant changes to the Human-AI Interaction paradigm (Morris et al., 2023). This level-based approach to defining AGI supports the coexistence of many prominent formulations – for example, Aguera y Arcas & Norvig’s definition (Agüera y Arcas & Norvig, 2023) would fall into the “Emerging AGI” category of our ontology, while OpenAI’s threshold of labor replacement (OpenAI, 2018) better matches “Virtuoso AGI.” Our “Competent AGI” level is probably the best catch-all for many existing definitions of AGI (e.g., the Legg (Legg, 2008), Shanahan (Shanahan, 2015), and Suleyman (Mustafa Suleyman and Michael Bhaskar, 2023) formulations). In the next section, we introduce a level-based ontology of AGI.
@@ 4 Levels of AGI
This taxonomy specifies the minimum performance over most tasks needed to achieve a given rating – e.g., a Competent AGI must have performance at least at the 50th percentile for skilled adult humans on most cognitive tasks, but may have Expert, Virtuoso, or even Superhuman performance on a subset of tasks. As an example of how individual systems may straddle different points in our taxonomy, we posit that as of this writing in September 2023, frontier language models (e.g., ChatGPT (OpenAI, 2023), Bard (Anil et al., 2023), Llama2 (Touvron et al., 2023), etc.) exhibit “Competent” performance levels for some tasks (e.g., short essay writing, simple coding), but are still at “Emerging” performance levels for most tasks (e.g., mathematical abilities, tasks involving factuality). Overall, current frontier language models would therefore be considered a Level 1 General AI (“Emerging AGI”) until the performance level increases for a broader set of tasks (at which point the Level 2 General AI, “Competent AGI,” criteria would be met). We suggest that documentation for frontier AI models, such as model cards (Mitchell et al., 2019), should detail this mixture of performance levels. This will help end-users, policymakers, and other stakeholders come to a shared, nuanced understanding of the likely uneven performance of systems progressing along the path to AGI.
@@
While this taxonomy rates systems according to their performance, systems that are capable of achieving a certain level of performance (e.g., against a given benchmark) may not match this level in practice when deployed. For instance, user interface limitations may reduce deployed performance. Consider DALLE-2 (Ramesh et al., 2022), which we estimate as a Level 3 Narrow AI (“Expert Narrow AI”) in our taxonomy. We estimate the “Expert” level of performance since DALLE-2 produces images of higher quality than most people are able to draw; however, the system has failure modes (e.g., drawing hands with incorrect numbers of digits, rendering nonsensical or illegible text) that prevent it from achieving a “Virtuoso” performance designation. While theoretically an “Expert” level system, in practice the system may only be “Competent,” because prompting interfaces are too complex for most end-users to elicit optimal performance (as evidenced by user studies (Zamfirescu-Pereira et al., 2023) and the existence of marketplaces (e.g., PromptBase ) in which skilled prompt engineers sell prompts). This observation emphasizes the importance of designing ecologically valid benchmarks (that approximate deployed rather than idealized performance), as well as the importance of considering the human-AI interaction paradigms.
@@ 6.1 Levels of AGI as a Framework for Risk Assessment
As we advance along our capability levels toward ASI, new risks are introduced, including misuse risks, alignment risks, and structural risks (Zwetsloot & Dafoe, 2019). For example, the “Expert AGI” level is likely to involve structural risks related to economic disruption and job displacement, as more and more industries reach the substitution threshold for machine intelligence in lieu of human labor. On the other hand, reaching “Expert AGI” likely alleviates some risks introduced by “Emerging AGI” and “Competent AGI,” such as the risk of incorrect task execution. The “Virtuoso AGI” and “ASI” levels are where many concerns relating to x-risk are most likely to emerge (e.g., an AI that can outperform its human operators on a broad range of tasks might deceive them to achieve a mis-specified goal, as in misalignment thought experiments (Christian, 2020)).
@@ 7 Conclusion
With these principles in mind, we introduced our Levels of AGI ontology, which offers a more nuanced way to define progress toward AGI by considering generality (either Narrow or General) in tandem with five levels of performance (Emerging, Competent, Expert, Virtuoso, and Superhuman). We reflected on how current AI systems and AGI definitions fit into this framing. Further, we discussed the implications of our principles for developing a living, ecologically valid AGI benchmark, and argue that such an endeavor, while sure to be challenging, is vital to engage with.

### v5

@@ 3 Defining AGI: Six Principles
6. Focus on the Path to AGI, not a Single Endpoint. Much as the adoption of a standard set of Levels of Driving Automation (SAE International, 2021) allowed for clear discussions of policy and progress relating to autonomous vehicles, we posit there is value in defining “Levels of AGI.” As we discuss in Section 5 and Section 6, we intend for each level of AGI to be associated with a clear set of metrics/benchmarks, as well as identified risks introduced at each level, and resultant changes to the Human-AI Interaction paradigm (Morris et al., 2023). This level-based approach to defining AGI supports the coexistence of many prominent formulations – for example, Aguera y Arcas & Norvig’s definition (Agüera y Arcas & Norvig, 2023) would fall into the “Emerging AGI” category of our ontology, while OpenAI’s threshold of labor replacement (OpenAI, 2018) better matches “Exceptional AGI.” Our “Competent AGI” level is probably the best catch-all for many existing definitions of AGI (e.g., the Legg (Legg, 2008), Shanahan (Shanahan, 2015), and Suleyman (Mustafa Suleyman and Michael Bhaskar, 2023) formulations). In the next section, we introduce a level-based ontology of AGI.
@@ 4 Levels of AGI
This taxonomy specifies the minimum performance over most tasks needed to achieve a given rating – e.g., a Competent AGI must have performance at least at the 50th percentile for skilled adult humans on most cognitive tasks, but may have Expert, Exceptional [^3] While Level 4 was originally called “Virtuoso AGI,” we now use the term “Exceptional AGI,” (Shah et al., 2025) which we believe better captures this capability level., or even Superhuman performance on a subset of tasks. As an example of how individual systems may straddle different points in our taxonomy, we posit that as of this writing in September 2023, frontier language models (e.g., ChatGPT (OpenAI, 2023), Bard (Anil et al., 2023), Llama2 (Touvron et al., 2023), etc.) exhibit “Competent” performance levels for some tasks (e.g., short essay writing, simple coding), but are still at “Emerging” performance levels for most tasks (e.g., mathematical abilities, tasks involving factuality). Overall, current frontier language models would therefore be considered a Level 1 General AI (“Emerging AGI”) until the performance level increases for a broader set of tasks (at which point the Level 2 General AI, “Competent AGI,” criteria would be met). We suggest that documentation for frontier AI models, such as model cards (Mitchell et al., 2019), should detail this mixture of performance levels. This will help end-users, policymakers, and other stakeholders come to a shared, nuanced understanding of the likely uneven performance of systems progressing along the path to AGI.
@@
While this taxonomy rates systems according to their performance, systems that are capable of achieving a certain level of performance (e.g., against a given benchmark) may not match this level in practice when deployed. For instance, user interface limitations may reduce deployed performance. Consider DALLE-2 (Ramesh et al., 2022), which we estimate as a Level 3 Narrow AI (“Expert Narrow AI”) in our taxonomy. We estimate the “Expert” level of performance since DALLE-2 produces images of higher quality than most people are able to draw; however, the system has failure modes (e.g., drawing hands with incorrect numbers of digits, rendering nonsensical or illegible text) that prevent it from achieving an “Exceptional” performance designation. While theoretically an “Expert” level system, in practice the system may only be “Competent,” because prompting interfaces are too complex for most end-users to elicit optimal performance (as evidenced by user studies (Zamfirescu-Pereira et al., 2023) and the existence of marketplaces (e.g., PromptBase ) in which skilled prompt engineers sell prompts). This observation emphasizes the importance of designing ecologically valid benchmarks (that approximate deployed rather than idealized performance), as well as the importance of considering the human-AI interaction paradigms.
@@ 6.1 Levels of AGI as a Framework for Risk Assessment
As we advance along our capability levels toward ASI, new risks are introduced, including misuse risks, alignment risks, and structural risks (Zwetsloot & Dafoe, 2019). For example, the “Expert AGI” level is likely to involve structural risks related to economic disruption and job displacement, as more and more industries reach the substitution threshold for machine intelligence in lieu of human labor. On the other hand, reaching “Expert AGI” likely alleviates some risks introduced by “Emerging AGI” and “Competent AGI,” such as the risk of incorrect task execution. The “Exceptional AGI” and “ASI” levels are where many concerns relating to x-risk are most likely to emerge (e.g., an AI that can outperform its human operators on a broad range of tasks might deceive them to achieve a mis-specified goal, as in misalignment thought experiments (Christian, 2020)).
@@ 7 Conclusion
With these principles in mind, we introduced our Levels of AGI ontology, which offers a more nuanced way to define progress toward AGI by considering generality (either Narrow or General) in tandem with five levels of performance (Emerging, Competent, Expert, Exceptional, and Superhuman). We reflected on how current AI systems and AGI definitions fit into this framing. Further, we discussed the implications of our principles for developing a living, ecologically valid AGI benchmark, and argue that such an endeavor, while sure to be challenging, is vital to engage with.